Paul had 30 marbles than Peter. After Peter gave Paul 15 more marbles, Paul had twice as many marbles as Peter. How many marbles did they have alltogether?
Answer by Paul(988)

Paul had 30 more marbles probably? or less?
If more so, then
Paul has x+30
Peter has x
Peter gave 15 marbles: -->x-15 so Paul had x+45 marbles now
Equation:
2(x-15)=x+45
2x-30=x+45
x=75
75+30=105
Hence, Peter has 75 marbles and Paul has 105 marbles.
Paul.